OpenStudy (anonymous):

Solve using the zero principle: (9s+5)(4s-12)=0

OpenStudy (anonymous):

first: 9s+5 = 0 9s = -5 s = -5/9 then: 4s - 12 = 0 4s = 12 s = 3

OpenStudy (anonymous):

zero principal means that if one part of a multiplication equation equals zero then the whole equation will so: 9s + 5 = 0 9s = -5 s = -9/5 4s-12=0 4s=12 s=3

OpenStudy (anonymous):

woops -5/9